OpenStudy (anonymous):

help :( not good in fractions 8 1/3 x 5 2/5 x 6 3/4 = ???

OpenStudy (anonymous):

81/3=27 52/5*63/4=13/5*63 so combing =27*63*13/5=22113/5=4422.6

OpenStudy (anonymous):

i mean 8(1/3) without the parenthesis

OpenStudy (anonymous):

i mean like this

OpenStudy (anonymous):

oh/...i see..wait.. 8/3*10/5*18/4=8/3*2*9/2=8/3*1*9=8*3=24 here * mean multiplication...